40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Thu Nov 7 14:39:49 2013, Jenne, Update, LSC, Lock Acquisition Game Plan LockAcquisition_7Nov2013.pdf
    Reply  Fri Nov 8 18:12:21 2013, Jenne, Update, LSC, PRFPMI: Not crossing any resonances 2f_resonances_Xarm_CARMsweep.pngAllModFreq_resonances_Xarm_CARMsweep.png
    Reply  Tue Nov 12 16:49:22 2013, Jenne, Update, LSC, Xend QPD and Whitening board pulled IMG_1262.JPGIMG_1260.JPGIMG_1264.JPGIMG_1276.JPG
       Reply  Wed Nov 20 17:05:15 2013, Jenne, Update, LSC, Xend QPD and Whitening board replaced 
       Reply  Wed Dec 18 00:26:15 2013, Jenne, Update, LSC, Xend QPD schematic investigation 
          Reply  Wed Dec 18 11:32:34 2013, rana, Update, LSC, Xend QPD schematic investigation 
    Reply  Wed Nov 13 01:35:40 2013, Jenne, Update, LSC, PRM motion causing trouble? xarmRIN_vs_CARM.pngASportOverlap_moveGaussianBeams.png
       Reply  Wed Nov 13 09:31:15 2013, Gabriele, Update, LSC, PRM motion causing trouble? 
          Reply  Wed Nov 13 19:22:58 2013, Jenne, Update, LSC, PRM motion correlated to intracavity power POPDC_sensors_30oct2013.pngAOMmodulated_POPDC_OSEM.pdf
             Reply  Thu Nov 14 02:50:43 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OL_TF_model.pngPRM_OLPIT_TF.pdfPRM_OLYAW_TF.pdf
                Reply  Mon Nov 18 21:02:54 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF PRM_OLG_NewOld.pngPRM_oplev_improvement.pdf
                   Reply  Mon Nov 18 21:40:24 2013, Koji, Update, LSC, PRM oplev measured and modeled TF 
                      Reply  Tue Nov 19 14:47:44 2013, Jenne, Update, LSC, PRM oplev measured and modeled TF 
             Reply  Thu Nov 14 02:55:26 2013, rana, Update, SUS, PRM motion correlated to intracavity power 
                Reply  Thu Nov 14 11:41:19 2013, Steve, Update, SUS, PRM sensors effected by IR IRon_off.png
                   Reply  Thu Nov 14 14:35:12 2013, Steve, Update, SUS, IR effect on MC and PRM sensors MC_PRM_IReffect.png
                      Reply  Fri Dec 6 10:03:07 2013, Steve, Update, SUS, IR effect on MC sensors only 10minPMCnotLocked.png6Dec2013.png
Message ID: 9383     Entry time: Thu Nov 14 02:55:26 2013     In reply to: 9378     Reply to this: 9384
Author: rana 
Type: Update 
Category: SUS 
Subject: PRM motion correlated to intracavity power 

 

Some more words about the ISS -> OSEM measurement:

The calibration of the OSEMs have been done so that these channels are each in units of microns. The SIDE channel has the lower noise floor because Valera increased the analog gain by 5x some time ago and compensated with lower digital gain.

The peak heights in the plot are:

UL   0.85

LL   0.78

UR   0.61

LR   0.45

S    0.27

So that tells us that the coupling is not uniform, but mostly coming in from the left side (which side is the the SIDE OSEM on?).

Jenne and I discussed what to do to mitigate this in the loops. Before we vent to fix the scattering (by putting some covers around the OSEMs perhaps), we want to try to tailor the OSEM damping loops to reduce their strength and increase the strength of the OL loops at the frequencies where we saw the bulk of the instability last time.

Jenne is optimizing OL loops now, and I'm working on OSEM tweaking. My aim is to lower the overall loop gains by ~3-5x and compensate that by putting in some low Q, resonant gain at the pendulum modes as we did for eLIGO. We did it here at the 40m several years ago, but had some troubles due to some resulting instability in the MC WFS loops.

 

In parallel, Steve is brainstorming some OSEM shields and I am asking around LIGO for some AC OSEM Satellite modules.

ELOG V3.1.3-